3. Given the initial firing angle 9 of a projectile, find the launch speed V , such that the projectile just hits the other side of the ravine a horizontal distance away of D
Answer #1

Expression for maximum range is given by

D= v^2 sin (2x)/g

v^2 = D* g / sin (2* theta)

v = [ D *g / sin (2*theta) ]

========

Comment in case any doubt.. Goodluck